# A Temporal-Spectral Fusion Transformer with Subject-specific Adapter for Enhancing RSVP-BCI Decoding

### v1

@@ 6.5.2 Feature Visualization
6.5.3 Subject Task-related Feature Extraction
In Fig. 7 (a), it can be observed that due to variations between different subjects and classes, EEG temporal signals exhibit very low similarity among subjects, with only weak similarities in the target data from different subjects. After feature extraction, the similarity of features among subjects increases (see Fig. 7 (b)). However, there are also notable similarities among features from different classes, which may increase FPR in classification. Figure 7 (c) illustrates that the similarity between features from different subjects within the same class is significantly enhanced compared to Figure 7 (b) and is much higher than the similarity between features from different classes. This improvement strengthens the TPR while reducing FPR in classification. These results demonstrate the model’s ability to effectively extract common task-related features across subjects through pre-training on data from various subjects.
6.5.4 Grad-CAM Visualization
Gradient-weighted Class Activation Mapping (Grad-CAM) is designed to reveal class-specific image regions from output feature maps using the class-specific gradient information while preserving the model structure selvaraju2017grad . We adopt Grad-CAM to visualize the attention of TSformer-SA to the target samples. Taking the first subject in Task plane as an example, the model is pre-trained on the other subjects in Task plane and the subject-specific adapter is fine-tuned on data from the first subject’s initial four blocks. Then, the target samples of the testing blocks are fed into the model, and two-view output features from the cross-view interaction module are collected to calculate the activation value for each target sample.
@@ 6.5.4 Grad-CAM Visualization
(a) Temporal view
(b) Spectral view
Figure 8: The Grad-CAM visualization results of TSformer-SA on the inputs of (a) temporal view and (b) spectral view of target samples. Because the inputs of two views are sliced and tokenized along the time dimension within the slice embedding layer, the activation value is one-dimensional. Shaded regions represent the activation value of that moment which is higher than the overall average.
Fig. 8 (a) and (b) show the activation values obtained by averaging the generated activation values of temporal and spectral views along the time dimension, respectively. Because we slice the inputs of two views along the time dimension and tokenize these slices into token sequences in the slice embedding layer, the activation value is one-dimensional. The shaded areas indicate moments with activation values higher than the overall average. For the temporal view, the high activation values are concentrated within two intervals: 260 ms to 324 ms and 360 ms to 696 ms, which correspond to the distinctive N200 and P300 components of EEG temporal signals in RSVP decoding, respectively. For the spectral view, the high activation values are mainly concentrated in the intervals of 256 ms to 500 ms and 720 ms to 776 ms, which are the discriminative parts of spectrogram images between target and nontarget samples. These outcomes demonstrate that our model can focus on the most discriminative regions of both EEG views for classification.
@@ 6.7 Computational Complexity in Preparation Procedure of BCI Systems
Before utilizing the BCI system, new subjects need to go through a preparation process which primarily consists of two parts: collecting labeled training data from new subjects and training individual models. In practical BCI applications, reducing preparation time is crucial for ensuring rapid deployment li2022tff ; wei2022erp ; wei2020reducing . The results in Section 6.6 indicate that TSformer-SA requires less training data from new subjects while achieving comparable performance to the comparison methods, which can shorten the time spent on collecting new subjects’ data during the preparation procedure of the RSVP-BCI system. In this part, we evaluate the time consumption of training an individual TSformer-SA during the preparation procedure.
@@
We calculate the number of trainable parameters and training time for TSformer-SA. For comparison, we list the results of EEGNet and PPNN, which possess fewer parameters than other CNN-based comparison methods. Additionally, we include EEG-conformer, which is the Transformer-based model with the fewest parameters. We take the Task people as an example and average the training time using the first block data from each subject. The experiments are performed on an Ubuntu server with the Intel Core i7-7700 2.10GHz CPU and NVIDIA GTX 2080 GPU.
As shown in Table 6, EEGNet has the fewest trainable parameters among the comparison methods. TSformer-SA possesses the fewest trainable parameters, accounting for only 54% of the parameter count in EEGNet. Notably, the number of trainable parameters in TSformer-SA is merely 1.5% of that in EEG-conformer. The average training time for TSformer-SA is 37.92 seconds, which differs by less than 1 second compared to EEGNet with the least training time. Thus, our proposed two-stage training strategy ensures that only the parameters of the subject-specific adapter are updated using new subject’s data during the preparation stage, which leads to a reduction in both computational costs and hardware requirements.
@@ 6.8 The Impact of Pre-training Data Sources
The results on the diagonal of the table represent the BA of TSformer-SA pre-trained and tested with the same RSVP task, while the results on the remaining positions are the BA of TSformer pre-trained and tested with different RSVP tasks. In each test task, the one-way repeated ANOVA reveals a significant main effect on different pre-training tasks (all test tasks: $p<0.001$). When the test task is Task plane and car respectively, there is no statistical difference between the BA of the TSformer-SA pre-trained with Task plane and car. but they are both significantly higher than the BA of TSformer-SA pre-trained with the Task people (all: $p<0.05$). When the test task switches to Task people, the BA of the TSformer pre-trained with Task people is significantly better than that of the TSformer pre-trained with Task plane and car (all: $p<0.05$). However, the BA fluctuations of TSformer-SA pre-trained on different tasks are only 1.32%, 1.30%, and 1.08% for Task plane, Task car, and Task people as test tasks, respectively. Meanwhile, TSformer-SA pre-trained on different tasks consistently outperforms the best comparison method in subject-dependent decoding on each test task (all tasks: $p<0.05$). For instance, when the Task plane serves as the test task, TSformer-SA pre-trained on the Task plane, car, and people all achieve significantly superior performance than HDCA which achieves the highest BA (88.05%) on the Task plane in subject-dependent decoding among all comparison methods (all: $p<0.05$).
@@
  • 1.
    The asterisks in the table indicate a significant difference between this result and the best result by paired t-tests (${}^{\star}$ $p<0.05$, ${}^{\star\star}$ $p<0.01$, ${}^{\star\star\star}$ $p<0.001$). The best results are highlighted in bold and the second-best results are denoted by underline.

### v2

@@ 6.7 Computational Complexity in Preparation Procedure of BCI Systems
Before utilizing the BCI system, new subjects need to go through a preparation process which primarily consists of two parts: collecting labeled training data from new subjects and training individual models. In practical BCI applications, reducing preparation time is crucial for ensuring rapid deployment [32, 66, 33]. The results in Section 6.6 indicate that TSformer-SA requires less training data from new subjects while achieving comparable performance to the comparison methods, which can shorten the time spent on collecting new subjects’ data during the preparation procedure of the RSVP-BCI system. In this part, we evaluate the time consumption of training an individual TSformer-SA during the preparation procedure.
@@
We calculate the number of trainable parameters and training time for TSformer-SA. For comparison, we list the results of EEGNet and PPNN, which possess fewer parameters than other CNN-based comparison methods. Additionally, we include EEG-conformer, which is the Transformer-based model with the fewest parameters. We take the Task people as an example since it has the highest number of samples per block among the three tasks. For each subject, the data from the first block are utilized for fine-tuning, where 80% of the data are randomly selected for training and the rest for validation. A training duration of 50 epochs per subject is conducted to determine the total training time, which is then averaged across subjects for the final results. The experiments are performed on an Ubuntu server with the Intel Core i7-7700 2.10GHz CPU and a NVIDIA GTX 2080 GPU.
As shown in Table 8, EEGNet has the fewest trainable parameters among the comparison methods. TSformer-SA possesses the fewest trainable parameters, accounting for only 54% of the parameter count in EEGNet. Notably, the number of trainable parameters in TSformer-SA is merely 1.5% of that in EEG-conformer. The average training time for TSformer-SA is 37.92 seconds, which differs by less than 1 second compared to EEGNet with the least training time. Conversely, the methods in [33] using adversarial transfer learning strategies takes 4.8 minutes of training time, which takes about eight times longer to that of TSformer-SA. Thus, our proposed two-stage training strategy ensures that only the parameters of the subject-specific adapter are updated using new subject’s data during the preparation stage, which leads to a reduction in both computational costs and hardware requirements.
@@ 6.8 The Impact of Pre-training Data Sources
The results on the diagonal of the table represent the BA of TSformer-SA pre-trained and tested with the same RSVP task, while the results on the remaining positions are the BA of TSformer pre-trained and tested with different RSVP tasks. In each test task, the one-way repeated ANOVA reveals a significant main effect on different pre-training tasks (all test tasks: $p<0.001$). When the test task is Task plane and car respectively, there is no statistical difference between the BA of the TSformer-SA pre-trained with Task plane and car, but they are both significantly higher than the BA of TSformer-SA pre-trained with the Task people (all: $p<0.05$). When the test task switches to Task people, the BA of the TSformer pre-trained with Task people is significantly better than that of the TSformer pre-trained with Task plane and car (all: $p<0.05$). However, the BA fluctuations of TSformer-SA pre-trained on different tasks are only 1.32%, 1.30%, and 1.08% for Task plane, Task car, and Task people as test tasks, respectively. Meanwhile, TSformer-SA pre-trained on different tasks consistently outperforms the best comparison method in subject-dependent decoding on each test task (all tasks: $p<0.05$). For instance, when the Task plane serves as the test task, TSformer-SA pre-trained on the Task plane, car, and people all achieve significantly superior performance than HDCA which achieves the highest BA (88.05%) on the Task plane in subject-dependent decoding among all comparison methods (all: $p<0.05$).
@@
  • The asterisks in the table indicate a significant difference between this result and the best result by paired t-tests ( $p<0.05$, ⋆⋆ $p<0.01$, ⋆⋆⋆ $p<0.001$). The best results are highlighted in bold and the second-best results are denoted by underline.
@@ 6.9 Sensitive Analysis of Wavelet Basis Function
6.10 Limitation and Future Work
Although our method achieves superior performance using EEG signals, it does not take advantage of multi-modal features such as eye movement signals and stimulus image features, which may further enhance RSVP decoding performance. In future research, we can explore a multi-modal learning approach that integrates EEG signals, eye movement data, and stimulus image characteristics to further enhance the decoding performance of RSVP tasks. Secondly, due to the limited dataset size, our model size is relatively small, which may restrict the decoding performance of our model across different subjects and datasets. In the future, we will try to expand the size of the dataset and then construct a pre-trained large model for RSVP tasks, which can improve the generalization ability of the decoding model across different subjects and datasets.